Justin is 4 years older than Ethan. In 9 years the sum of their ages will be 30. Find the age of each boy now.
Justin is ? years
Ethan is ? years

Answer by checkley79(3341)

Justin=x+4
Ethan=x
(x+4)+9+x+9=30
2x=30-22
2x=8
x=8/2
x=4 the age of Ethan.
4+4=8 the age of Justin.
Proof:
8+9+4+9=30
330=30